OpenStudy (anonymous):

FAN & MEDAL !! a sample of CH4 is burned in an excess of O2 to give 4.2 moles of H2O how many moles of CH4 were in the original sample ? I got 2.1 as an answer but don't know if that is correct or not and my paper is graded

OpenStudy (anonymous):

yes 2.1 is correct

OpenStudy (abb0t):

Well, let's see. Let's start off by writing down the balanced chemical equation for combustion of methane, \(\sf \color{}{CH_4}\): \(\sf \color{red}{CH_4 + 2~O_2 \rightarrow} \color{blue}{CO_2+ 2~H_2O}\) so, your work should be as followed: \(\sf 4.2~mol~H_2O \times \frac{1~mol~CH_4}{2~mol~O_2}\) and as you can clearly see, the two units cancel out nicely! \(\sf 4.2\cancel{~mol~H_2O} \times \frac{1~mol~CH_4}{2~ \cancel{~mol~O_2}}\) = \(\sf \frac{4.2}{2}~mol~CH_4\) = \(\sf \color{red}{2.1} ~mol~of~CH_4\)

OpenStudy (abb0t):

Sorry, I wrote down 2 moles of \(\sf O_2\) when instead I mean't to write \(\sf \LARGE \color{purple}{H_2O}\). My mistake.